OpenStudy (anonymous):

Multiply (3)(−4)(2). −48 −24 24 48

OpenStudy (anonymous):

3*-4*2 = -12*2 = -24

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Multiply (−4)(−2)(−5). −40 −8 8 40

OpenStudy (anonymous):

(-4)*(-2)*(-5) = -40 You have to know that \[+*+ = +\] \[-*- = +\] \[+*- = -\] \[-*+ = -\] Than you can answer anykind of these questions